The difference of triple a number and 1/2 is equal to the sum of the number and 2/3. Asked to find the number
Answer by stanbon(75887)

The difference of triple a number and 1/2 is equal to the
sum of the number and 2/3.
------------------
3x - (1/2) = x + (2/3)
------
2x = (2/3)+(1/2)
----
2x = (4+3)/6
x = 7/6
==============
Cheers,
Stan H.
=================